OpenStudy (anonymous):

so how do you draw the angle abc meassures 65

OpenStudy (anonymous):

ok Q. Find the m angle ADB in the given ΔABC, if AD is the angle bisector of angle BAC and if m angle ABC = 75 and m angle BCA = 65 Choices are: a. 65 b. 85 c. 75 d. 55 Like Dislike Answers Answer 1 m angle ABC = 75 and m angle BCA = 65 [Given.] The sum of the measures of the angles of a triangle is 180. [Triangle -Angle sum theorem.] So, m angle BAC + m angle ABC + m angle BCA = 180 m angle BAC + 75 + 65 = 180 [Substitute given measures.] m angle BAC = 180 - 140 = 40 AD is the angle bisector of angle BAC [Given.] So, m angle BAD = m∠BAC/2 = 40/2 = 20 The sum of the measures of the angles of a triangle is 180. [Triangle angle sum theorem.] m angle BAD + m angle ABD + m angle ADB = 180 20 + 75 + m angle ADB = 180 [Substitute given measures.] m angle ADB = 180 - 95 = 85 Thus, correct answer is option 'B'.

OpenStudy (anonymous):

Q. Find the m angle ADB in the given ΔABC, if AD is the angle bisector of angle BAC and if m angle ABC = 75 and m angle BCA = 65 Choices are: a. 65 b. 85 c. 75 d. 55 m angle ABC = 75 and m angle BCA = 65 [Given.] The sum of the measures of the angles of a triangle is 180. [Triangle -Angle sum theorem.] So, m angle BAC + m angle ABC + m angle BCA = 180 m angle BAC + 75 + 65 = 180 [Substitute given measures.] m angle BAC = 180 - 140 = 40 AD is the angle bisector of angle BAC [Given.] So, m angle BAD = m∠BAC/2 = 40/2 = 20 The sum of the measures of the angles of a triangle is 180. [Triangle angle sum theorem.] m angle BAD + m angle ABD + m angle ADB = 180 20 + 75 + m angle ADB = 180 [Substitute given measures.] m angle ADB = 180 - 95 = 85 Thus, correct answer is option 'B'.